a car of mas 1000kg moving with a speed 18km/h on a smooth road and colliding with a horizontally mounted spring of spring constant 6.25x10 cube. wht is thw max compression of the spring?
 

Grade:11

1 Answers

Manas Shukla
7 years ago
Just equate the energy.
18 km/hr = 5 m/s
Energy of car = \frac{1}{2}mv^{2} = 25000/2
Energy of spring = \frac{1}{2}kx^{2}
Equating we get 
x = v (m/k)^{\frac{1}{2}} = 2 m